WHAT IS THE PERIMETER OF A RECTANGLE IF THE LENGHT IS 7/10 IN. AND THE WIDTH IS 1/10IN.
Answer by jojo14344(1512)

Given,
L= 7/10"
W=1/10"
P= 2(L+W)
=2(7/10 + 1/10)
=2(8/10)=2(4/5) = 1-3/5 in. or
P=1.6 inches
Thank you,
Jojo